# Subvarieties of pointed Abelian $\ell$-groups

### v1

@@ 4 Characterization of varieties generated by a single finitely generated totally ordered $\ell$-group
we get ${\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})\in\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}).$ This proves the claim. ∎
@@
Let ${\boldsymbol{\mathit{A}}}_{a}$ be a finitely generated totally ordered Abelian $\ell$-group.
@@
By Theorem 3.10 we can assume without loss of generality that ${\boldsymbol{\mathit{A}}}_{a}$ is strongly pointed. Let us assume $\mathbf{rank}({\boldsymbol{\mathit{A}}}_{a})=\infty$. Since $\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})$ is a maximal convex $\ell$-subgroup of ${\boldsymbol{\mathit{A}}}_{a}$, the Abelian $\ell$-group ${\boldsymbol{\mathit{A}}}_{a}/\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})$ has to be p-simple and thus by Theorem 2.8 ${\boldsymbol{\mathit{A}}}_{a}/\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})\in\mathbf{IS}({\boldsymbol{\mathit{R}}}_{b})$ for some $0<b\in\mathbb{R}$. Let us without loss of generality assume ${\boldsymbol{\mathit{A}}}_{a}/\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})\in\mathbf{S}({\boldsymbol{\mathit{R}}}_{1})$. By Lemma 2.9 we obtain ${\boldsymbol{\mathit{A}}}_{a}/\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})\cong{\boldsymbol{\mathit{Z}}}_{n}$ for some $n\in\mathbb{Z}$ or the universe of ${\boldsymbol{\mathit{A}}}_{a}/\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})$ is a dense subset of $\mathbb{R}$. Since $\mathbf{rank}({\boldsymbol{\mathit{A}}}_{a})=\infty$, the universe of ${\boldsymbol{\mathit{A}}}_{a}/\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})$ is dense in $\mathbb{R}$. Thus by Lemma 4.3 we get $\mathbf{HSP}({\boldsymbol{\mathit{A}}}_{a}/\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a}))=\mathbf{HSP}({\boldsymbol{\mathit{R}}}_{1})$. Since ${\boldsymbol{\mathit{A}}}_{a}/\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})\in\mathbf{H}({\boldsymbol{\mathit{A}}}_{a})$ and by Lemma 3.7 ${\boldsymbol{\mathit{A}}}_{a}\in\mathbf{HSP}({\boldsymbol{\mathit{R}}}_{1})$ we obtain $\mathbf{HSP}({\boldsymbol{\mathit{R}}}_{1})=\mathbf{HSP}({\boldsymbol{\mathit{A}}}_{a})$.
@@
All together we obtained the following characterization.
@@
Let ${\boldsymbol{\mathit{A}}}_{a}$ be a finitely generated totally ordered pointed Abelian $\ell$-group. Then the following holds.
@@
    $\mathbf{HSP}({\boldsymbol{\mathit{A}}}_{a})=\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})$ iff ${\boldsymbol{\mathit{A}}}_{a}$ is non-p-simple and $\mathbf{rank}({\boldsymbol{\mathit{A}}}_{a})=n$.
@@ 5 Axiomatization of subvarieties of $p\mathbb{AL}$
In this section we will introduce several equations which we later use for axiomatizations of subvarieties of $p\mathbb{AL}$. We will introduce three important equations. It should be noted that all our equations are valid in all negatively pointed $\ell$-groups.
@@
Let ${\boldsymbol{\mathit{A}}}_{a}$ be a totally ordered pointed Abelian $\ell$-group and $n\geq 0$. Let us consider the following equation.
@@
By Theorem 3.10 we can assume ${\boldsymbol{\mathit{A}}}_{a}$ is strongly pointed. Thus by Lemma 4.7 we need to check that
@@
Let $e$ be a fixed evaluation on ${\boldsymbol{\mathit{A}}}_{a}$. We have ${\boldsymbol{\mathit{A}}}_{a}\vDash_{e}\mathrm{(\text{s-rank}_{n})}$ iff ${\boldsymbol{\mathit{A}}}_{a}\vDash_{e}e(x)\leq 0$ or ${\boldsymbol{\mathit{A}}}_{a}\vDash_{e}n\cdot e(x)\geq a$. Therefore ${\boldsymbol{\mathit{A}}}_{a}\vDash\mathrm{(\text{s-rank}_{n})}$ iff ${\boldsymbol{\mathit{A}}}_{a}\vDash n\cdot e(x)\geq a$ for all evaluations $e$, such that $e(x)>0$.
@@
    For $a>0$ and ${\boldsymbol{\mathit{A}}}_{a}$ non-p-simple we have $\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})\neq 0$ and by Lemma 3.2 there is an isomorphism $\iota:{\boldsymbol{\mathit{A}}}_{a}/\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})\overrightarrow{\times}\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})\rightarrow{\boldsymbol{\mathit{A}}}_{a}$. Pick $0<v\in\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})$ and we consider evaluation $e$ on $\boldsymbol{\mathit{A}}$}, where $e(x)=\iota(\langle{0,v}\rangle)$. Since $n\cdot v\in\mathbf{p\textnormal{-}rad}({\boldsymbol{\mathit{A}}}_{a})$, we have $n\cdot e(x)=\iota(\langle{0,n\cdot v}\rangle)<a$, thus showing ${\boldsymbol{\mathit{A}}}_{a}\nvDash\mathrm{(\text{s-rank}_{n})}$ for non-p-simple ${\boldsymbol{\mathit{A}}}_{a}$ with $a>0$.
Let us note that $\mathrm{(\text{s-rank}_{0})}$ is equivalent to the equation $f\leq 0$. For $n\in\mathbb{Z}$ we denote $\operatorname{div}(z)$ to be a set of all divisors of $z$.
@@
Since ${\boldsymbol{\mathit{Z}}}_{m}\in\mathbf{H}({\boldsymbol{\mathit{Z}}}_{m}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})$ it is enough to check that ${\boldsymbol{\mathit{Z}}}_{m}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\vDash\mathrm{(div_{p,n})}$ if $m\leq 0$ or $p\notin\operatorname{div}(m)$ and that ${\boldsymbol{\mathit{Z}}}_{m}\nvDash\mathrm{(div_{p,n})}$ if $0<m$ and $p\in\operatorname{div}(m)$.
@@
    Now assume $0<m\leq n,p\notin\operatorname{div}(m)$ and $e(x)=\langle{a_{1},a_{2}}\rangle$ and $a_{1}\geq 0$. Since $p\notin\operatorname{div}(m)$ we get $p\cdot a_{1}\neq m$ and thus $|p\cdot a_{1}-m|\geq 1$. Consequently,
@@
Now assume ${\boldsymbol{\mathit{A}}}_{a}$ is an arbitrary finitely generated subdirectly irreducible positively pointed Abelian $\ell$-group, such that ${\boldsymbol{\mathit{A}}}_{a}\vDash\mathrm{(\text{s-rank}_{n})}$ and ${\boldsymbol{\mathit{A}}}_{a}\vDash\mathrm{(\text{div}_{p,{n}})}$ for all $p\notin\operatorname{div}(n)$. We will show ${\boldsymbol{\mathit{A}}}_{a}\in\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{n})$.
By Theorem 2.5 ${\boldsymbol{\mathit{A}}}_{a}$ is totally ordered. By Theorem 3.10 we can assume without loss of generality that ${\boldsymbol{\mathit{A}}}_{a}$ is strongly pointed. By Lemma 5.1 we obtain ${\boldsymbol{\mathit{A}}}_{a}$ has to be p-simple with rank $n$ and thus by Lemma 4.7 ${\boldsymbol{\mathit{A}}}_{a}\cong{\boldsymbol{\mathit{Z}}}_{k}$ for some $k\leq n$. By Lemma 5.2 we obtain $k$ divides $n$. That means ${\boldsymbol{\mathit{A}}}_{a}\cong{\boldsymbol{\mathit{Z}}}_{k}\in\mathbf{IS}(Z_{n})$ and thus ${\boldsymbol{\mathit{A}}}_{a}\in\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{n})$. ∎
Since Theorem 5.7 also provides an axiomatization for the variety generated by ${\boldsymbol{\mathit{Z}}}_{n}$, it might seem that this would render Theorem 5.3 redundant. However, this is not the case. The varieties generated by ${\boldsymbol{\mathit{Z}}}_{n}$ are structurally simpler than the general case, as they are not generated by any $\ell$-groups of the form ${\boldsymbol{\mathit{Z}}}_{m}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}$. Consequently, Theorem 5.3 provides a much simpler and more direct axiomatization than the one required by the general framework of Theorem 5.7. This is particularly appealing because the varieties generated by ${\boldsymbol{\mathit{Z}}}_{n}$ correspond via the Mundici functor to the varieties generated by single finite Łukasiewicz chains, which are among the most fundamental structures in the theory of MV-algebras.
@@
$((2n+1)\cdot x-2\cdot\mathtt{f})\lor(\mathtt{f}-(2n+2)\cdot x)\lor-x\geq 0.$ ($rank_{n}$)
@@
    First we show ${\boldsymbol{\mathit{Z}}}_{m}\nvDash\mathrm{(\text{rank}_{n})}$ for $n<m$. Since ${\boldsymbol{\mathit{Z}}}_{m}$ is Archimedean, there is a maximal $k\geq 1$ such that $k\cdot(2n+1)<2m$. Let us note, that from the maximality of $k$ it follows $2k\cdot(2n+1)\geq 2m$ and thus $k\cdot(2n+1)\geq m$. Let $e$ be an evaluation on ${\boldsymbol{\mathit{Z}}}_{m}$ such that $e(x)=k$. Now we have
@@
    by definition of $k$. Moreover, we have
@@
    This shows that indeed ${\boldsymbol{\mathit{Z}}}_{m}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\vDash\mathrm{(\text{rank}_{n})}$ for $m\leq n$.
@@
Let $\boldsymbol{\mathit{A}}$} be totally ordered Abelian $\ell$-group and $0\leq p\leq n$. Let us consider the following equation.
$((n+1)\cdot((p\cdot x-\mathtt{f})\lor(\mathtt{f}-p\cdot x))-\mathtt{f})\lor(n\cdot y-\mathtt{f})\lor(-y)\geq 0$ (mixp,n)
@@
Any proper subvariety of $p\mathbb{AL}^{+}$ is of the form
$V_{I,J}=\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{i},{\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\mid i\in I,j\in J)$
for some finite sets $J\subseteq I\subsetneq\mathbb{N}$, where both $I$ and $J$ are closed under divisors.
Moreover, $V_{I,J}$ is axiomatized by the following set of equations:
@@
First we show that any proper subvariety of $p\mathbb{AL}^{+}$ is equal to $V_{I,J}$ for some finite sets $J\subseteq I$.
Let $\mathbb{K}$}} be an arbitrary subvariety of $p\mathbb{AL}^{+}$. Let us denote $I=\{i\mid{\boldsymbol{\mathit{Z}}}_{i}\in{{\mathbb{K}}}\}$ and $J=\{j\mid{\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\in{{\mathbb{K}}}\}$. We show ${{\mathbb{K}}}=V_{I,J}$. Since $\mathbb{K}$}} is a proper subvariety of $p\mathbb{AL}^{+}$, by Lemma 3.7 we know that ${\boldsymbol{\mathit{R}}}_{1}\notin{{\mathbb{K}}}$. By Theorem 4.10 we have
@@
Clearly, both $I$ and $J$ are closed under divisors, since for any $d,k\in\mathbb{N}\setminus\{0\}$ such that $d\in\operatorname{div}(k)$ we have ${\boldsymbol{\mathit{Z}}}_{d}\in\mathbf{IS}({\boldsymbol{\mathit{Z}}}_{k})$ and ${\boldsymbol{\mathit{Z}}}_{d}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\in\mathbf{IS}({\boldsymbol{\mathit{Z}}}_{k}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})$. Since for each $j\in J$ we have ${\boldsymbol{\mathit{Z}}}_{j}\in\mathbf{H}({\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})$, we get $J\subseteq I$. The set $I$ (and consequently $J$) has to be finite, otherwise we would get by Lemma 4.3 that ${\boldsymbol{\mathit{R}}}_{1}\in{{\mathbb{K}}}$. This proves ${{\mathbb{K}}}=V_{I,J}$ for some finite sets $J\subseteq I$.
@@
Now we have to show the axiomatization of $V_{I,J}$. First, we will argue that all the equations from $S_{I,J}$ hold in $V_{I,J}$. Since $n=\max I\geq\max J$, by Lemma 5.4 we know that ${\boldsymbol{\mathit{Z}}}_{i}\vDash\mathrm{(rank_{n})}$ for all $i\in I$ and ${\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\vDash\mathrm{(rank_{n})}$ for all $j\in J$. Thus $V_{I,J}\vDash\mathrm{(rank_{n})}$.
@@
Let us fix $p\notin I$. Since $I$ is closed under divisors, we get that $i\notin\operatorname{div}(p)$ for all $i\in I$. Using $J\subseteq I$ and by Lemma 5.2 we obtain that ${\boldsymbol{\mathit{Z}}}_{i}\vDash\mathrm{(div_{p,n})}$ for each $i\in I$ and ${\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\vDash\mathrm{(div_{p,n})}$ for each $j\in J$. Therefore, $V_{I,J}\vDash\mathrm{(div_{p,n})}$ for each $p\notin I$.

### v2

@@ 3 The lexicographic decompositions
we get ${\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}\mathbf{val}({\boldsymbol{\mathit{A}}}_{a})\in\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}).$ This proves the claim. ∎
@@
Let ${\boldsymbol{\mathit{A}}}_{a}$ be a finitely generated totally ordered pointed Abelian $\ell$-group.
@@
By Theorem 3.10 we can assume without loss of generality that ${\boldsymbol{\mathit{A}}}_{a}$ is strongly pointed. Let us assume $\mathbf{rank}({\boldsymbol{\mathit{A}}}_{a})=\infty$. Since $\mathbf{val}({\boldsymbol{\mathit{A}}}_{a})$ is a maximal convex $\ell$-subgroup of ${\boldsymbol{\mathit{A}}}_{a}$, the Abelian $\ell$-group ${\boldsymbol{\mathit{A}}}_{a}/\mathbf{val}({\boldsymbol{\mathit{A}}}_{a})$ must be p-simple and thus by Theorem 2.9 ${\boldsymbol{\mathit{A}}}_{a}/\mathbf{val}({\boldsymbol{\mathit{A}}}_{a})\in\mathbf{IS}({\boldsymbol{\mathit{R}}}_{b})$ for some $0<b\in\mathbb{R}$. Let us without loss of generality assume ${\boldsymbol{\mathit{A}}}_{a}/\mathbf{val}({\boldsymbol{\mathit{A}}}_{a})\in\mathbf{S}({\boldsymbol{\mathit{R}}}_{1})$. By Lemma 2.10 we obtain ${\boldsymbol{\mathit{A}}}_{a}/\mathbf{val}({\boldsymbol{\mathit{A}}}_{a})\cong{\boldsymbol{\mathit{Z}}}_{n}$ for some $n\in\mathbb{Z}$ or the universe of ${\boldsymbol{\mathit{A}}}_{a}/\mathbf{val}({\boldsymbol{\mathit{A}}}_{a})$ is a dense subset of $\mathbb{R}$. Since $\mathbf{rank}({\boldsymbol{\mathit{A}}}_{a})=\infty$, the universe of ${\boldsymbol{\mathit{A}}}_{a}/\mathbf{val}({\boldsymbol{\mathit{A}}}_{a})$ is dense in $\mathbb{R}$. Thus by Lemma 4.3 we get $\mathbf{HSP}({\boldsymbol{\mathit{A}}}_{a}/\mathbf{val}({\boldsymbol{\mathit{A}}}_{a}))=\mathbf{HSP}({\boldsymbol{\mathit{R}}}_{1})$. Since ${\boldsymbol{\mathit{A}}}_{a}/\mathbf{val}({\boldsymbol{\mathit{A}}}_{a})\in\mathbf{H}({\boldsymbol{\mathit{A}}}_{a})$ and by Lemma 3.7 ${\boldsymbol{\mathit{A}}}_{a}\in\mathbf{HSP}({\boldsymbol{\mathit{R}}}_{1})$ we obtain $\mathbf{HSP}({\boldsymbol{\mathit{R}}}_{1})=\mathbf{HSP}({\boldsymbol{\mathit{A}}}_{a})$.
@@
Altogether, we obtained the following characterization.
@@
Let ${\boldsymbol{\mathit{A}}}_{a}$ be a finitely generated non-trivial totally ordered pointed Abelian $\ell$-group. Then the following holds.
@@
    $\mathbf{HSP}({\boldsymbol{\mathit{A}}}_{a})=\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{n}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})$ iff ${\boldsymbol{\mathit{A}}}_{a}$ is not p-simple and $\mathbf{rank}({\boldsymbol{\mathit{A}}}_{a})=n$.
@@ 5 Axiomatization of subvarieties of $p\mathbb{AL}$
In this section we focus on axiomatizing all the subvarieties of $p\mathbb{AL}^{+}$. To achieve this we define the following equations: $\mathrm{(\text{s-rank}_{n})}$$\mathrm{(\text{rank}_{n})}$, $\mathrm{(\text{div}_{p,n})}$ and $\mathrm{(\text{mix}_{p,n})}$. Observe that all these four equations are valid for all parameters in all negatively pointed Abelian $\ell$-groups. Consequently, we introduce another four dual equations $\mathrm{(\text{s-rank}^{d}_{n})}$$\mathrm{(\text{rank}^{d}_{n})}$, $\mathrm{(\text{div}^{d}_{p,n})}$ and $\mathrm{(\text{mix}^{d}_{p,n})}$, which we use to characterize all subvarieties of $p\mathbb{AL}^{-}$. Finally, we combine these strategies to provide a characterization of all subvarieties of $p\mathbb{AL}$.
@@
Let ${\boldsymbol{\mathit{A}}}_{a}$ be a finitely generated totally ordered pointed Abelian $\ell$-group and $n\geq 0$. Let us consider the following equation.
@@
First, if $a=0$, the equation is satisfied by Corollary 3.6. Now, we can assume $a\neq 0$. By Theorem 3.10 we can assume ${\boldsymbol{\mathit{A}}}_{a}$ is strongly pointed. Thus by Lemma 4.7 we need to check that
@@
Let $e$ be a fixed evaluation on ${\boldsymbol{\mathit{A}}}_{a}$. We have ${\boldsymbol{\mathit{A}}}_{a}\vDash_{e}\mathrm{(\text{s-rank}_{n})}$ iff ${\boldsymbol{\mathit{A}}}_{a}\vDash_{e}x\leq 0$ or ${\boldsymbol{\mathit{A}}}_{a}\vDash_{e}n\cdot x\geq a$. Therefore ${\boldsymbol{\mathit{A}}}_{a}\vDash\mathrm{(\text{s-rank}_{n})}$ iff ${\boldsymbol{\mathit{A}}}_{a}\vDash_{e}n\cdot x\geq a$ for all evaluations $e$, such that $e(x)>0$.
@@
    For $a>0$ and ${\boldsymbol{\mathit{A}}}_{a}$ not p-simple we have $\mathbf{val}({\boldsymbol{\mathit{A}}}_{a})\neq 0$ and by Lemma 3.2 there is an isomorphism $\iota:{\boldsymbol{\mathit{A}}}_{a}/\mathbf{val}({\boldsymbol{\mathit{A}}}_{a})\overrightarrow{\times}\mathbf{val}({\boldsymbol{\mathit{A}}}_{a})\rightarrow{\boldsymbol{\mathit{A}}}_{a}$. Pick $0<v\in\mathbf{val}({\boldsymbol{\mathit{A}}}_{a})$ and we consider an evaluation $e$ on $\boldsymbol{\mathit{A}}$}, where $e(x)=\iota(\langle{0,v}\rangle)$. Since $n\cdot v\in\mathbf{val}({\boldsymbol{\mathit{A}}}_{a})$, we have $n\cdot e(x)=\iota(\langle{0,n\cdot v}\rangle)<a$, thus showing ${\boldsymbol{\mathit{A}}}_{a}\nvDash\mathrm{(\text{s-rank}_{n})}$ for not p-simple ${\boldsymbol{\mathit{A}}}_{a}$ with $a>0$. ∎
Let us note that $\mathrm{(\text{s-rank}_{0})}$ is equivalent to the equation $f\leq 0$. For $m\in\mathbb{Z}\setminus\{0\}$ we denote $\operatorname{div}(m)$ to be the set of all divisors of $m$.
@@
Since ${\boldsymbol{\mathit{Z}}}_{m}\in\mathbf{H}({\boldsymbol{\mathit{Z}}}_{m}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})$, it is enough to check that ${\boldsymbol{\mathit{Z}}}_{m}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\vDash\mathrm{(div_{p,n})}$ if $m\leq 0$ or $p\notin\operatorname{div}(m)$ and that ${\boldsymbol{\mathit{Z}}}_{m}\nvDash\mathrm{(div_{p,n})}$ if $0<m$ and $p\in\operatorname{div}(m)$.
@@
    Now assume $0<m\leq n,p\notin\operatorname{div}(m)$, $e(x)=\langle{a_{1},a_{2}}\rangle$ and $a_{1}\geq 0$. Since $p\notin\operatorname{div}(m)$ we get $p\cdot a_{1}\neq m$, and thus $|p\cdot a_{1}-m|\geq 1$. Consequently,
@@
Now assume there is a finitely generated ${\boldsymbol{\mathit{A}}}_{a}\in p\mathbb{AL}^{+}_{\mathrm{FSI}}$, such that ${\boldsymbol{\mathit{A}}}_{a}\vDash\mathrm{(\text{s-rank}_{n})}$ and ${\boldsymbol{\mathit{A}}}_{a}\vDash\mathrm{(\text{div}_{p,{n}})}$ for all $p\notin\operatorname{div}(n)$. We will show ${\boldsymbol{\mathit{A}}}_{a}\in\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{n})$.
By Theorem 2.6, ${\boldsymbol{\mathit{A}}}_{a}$ is totally ordered. If $a=0$ we obtain the desired result by applying Corollary 3.6. Let us now assume that $a\neq 0$. By Theorem 3.10 we can assume without loss of generality that ${\boldsymbol{\mathit{A}}}_{a}$ is strongly pointed. By Lemma 5.1 we obtain ${\boldsymbol{\mathit{A}}}_{a}$ must be p-simple with rank less or equal to $n$ and thus by Lemma 4.7 ${\boldsymbol{\mathit{A}}}_{a}\cong{\boldsymbol{\mathit{Z}}}_{k}$ for some $k\leq n$. By Lemma 5.2 we obtain $k$ divides $n$. That means ${\boldsymbol{\mathit{A}}}_{a}\cong{\boldsymbol{\mathit{Z}}}_{k}\in\mathbf{IS}({\boldsymbol{\mathit{Z}}}_{n})$ and thus ${\boldsymbol{\mathit{A}}}_{a}\in\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{n})$. ∎
Since Theorem 5.7 also provides an axiomatization for the variety generated by ${\boldsymbol{\mathit{Z}}}_{n}$, it might seem that this would render Theorem 5.3 redundant. However, this is not the case. The varieties generated by ${\boldsymbol{\mathit{Z}}}_{n}$ are structurally simpler than the general case, as they are not generated by any $\ell$-groups of the form ${\boldsymbol{\mathit{Z}}}_{m}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}$. Consequently, Theorem 5.3 provides a much simpler and more direct axiomatization than the one required by the general framework of Theorem 5.7. This is particularly appealing because the varieties generated by ${\boldsymbol{\mathit{Z}}}_{n}$ correspond via the Mundici functor to the varieties generated by single finite simple MV-algebra, which are among the most fundamental structures in the theory of MV-algebras.
@@
$((2n+1)\cdot x-2\cdot\mathtt{f})\lor(\mathtt{f}-(2n+2)\cdot x)\lor-x\geq 0.$ (rankn)
@@
    First we show ${\boldsymbol{\mathit{Z}}}_{m}\nvDash\mathrm{(\text{rank}_{n})}$ for $n<m$. Since ${\boldsymbol{\mathit{Z}}}_{m}$ is Archimedean, there is a maximal $k\geq 1$ such that $k\cdot(2n+1)<2m$. Let us note that from the maximality of $k$ it follows $2k\cdot(2n+1)\geq 2m$ and thus $k\cdot(2n+1)\geq m$. Let $e$ be an evaluation on ${\boldsymbol{\mathit{Z}}}_{m}$ such that $e(x)=k$. Now we have
@@
    by the definition of $k$. Moreover, we have
@@
    This shows that indeed ${\boldsymbol{\mathit{Z}}}_{m}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\vDash\mathrm{(\text{rank}_{n})}$ for $m\leq n$. ∎
@@
Let $\boldsymbol{\mathit{A}}$} be a finitely generated totally ordered pointed Abelian $\ell$-group and $0\leq p\leq n$. Let us consider the following equation.
$((n+1)\cdot((p\cdot x-\mathtt{f})\lor(\mathtt{f}-p\cdot x))-\mathtt{f})\lor(-x)\lor(n\cdot y-\mathtt{f})\lor(-y)\geq 0$ (mixp,n)
@@
Any non-trivial proper subvariety of $p\mathbb{AL}^{+}$ is of the form
$\mathcal{V}_{I,J}=\mathbf{HSP}({\boldsymbol{\mathit{Z}}}_{i},{\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\mid i\in I,j\in J)$
for some finite sets $\{0\}\subseteq J\subseteq I\subsetneq\mathbb{N}$, where both $I\setminus\{0\}$ and $J\setminus\{0\}$ are closed under divisors.
Moreover, $\mathcal{V}_{I,J}$ is axiomatized by the following set of equations:
@@
First we show that any proper subvariety of $p\mathbb{AL}^{+}$ is equal to $\mathcal{V}_{I,J}$ for some finite sets $\{0\}\subseteq J\subseteq I$. Let $\mathcal{V}$ be an arbitrary subvariety of $p\mathbb{AL}^{+}$. Let us denote $I=\{i\mid{\boldsymbol{\mathit{Z}}}_{i}\in\mathcal{V}\}$ and $J=\{j\mid{\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\in\mathcal{V}\}$. By Corollary 3.6 we have $\{0\}\in I\cap J$. We show $\mathcal{V}=\mathcal{V}_{I,J}$. Since $\mathcal{V}$ is a proper subvariety of $p\mathbb{AL}^{+}$, by Lemma 3.7 we know that ${\boldsymbol{\mathit{R}}}_{1}\notin\mathcal{V}$. By Theorem 4.10 we have
@@
Clearly, both $I\setminus\{0\}$ and $J\setminus\{0\}$ are closed under divisors, since for any $d,k\in\mathbb{N}\setminus\{0\}$ such that $d\in\operatorname{div}(k)$ we have ${\boldsymbol{\mathit{Z}}}_{d}\in\mathbf{IS}({\boldsymbol{\mathit{Z}}}_{k})$ and ${\boldsymbol{\mathit{Z}}}_{d}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\in\mathbf{IS}({\boldsymbol{\mathit{Z}}}_{k}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})$. Since for each $j\in J$ we have ${\boldsymbol{\mathit{Z}}}_{j}\in\mathbf{H}({\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0})$, we get $J\subseteq I$. The set $I$ (and consequently $J$) must be finite, otherwise we would get by Lemma 4.3 that ${\boldsymbol{\mathit{R}}}_{1}\in\mathcal{V}$. This proves $\mathcal{V}=\mathcal{V}_{I,J}$ for some finite sets $J\subseteq I$.
@@
Now we have to show the axiomatization of $\mathcal{V}_{I,J}$. First, we will argue that all the equations from $S_{I,J}$ hold in $\mathcal{V}_{I,J}$. Since $n=\max I\geq\max J$, by Lemma 5.4 we know that ${\boldsymbol{\mathit{Z}}}_{i}\vDash\mathrm{(rank_{n})}$ for all $i\in I$ and ${\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\vDash\mathrm{(rank_{n})}$ for all $j\in J$. Thus $\mathcal{V}_{I,J}\vDash\mathrm{(rank_{n})}$.
@@
Let us fix $p\notin I$. Since $I$ is closed under divisors, we get that $p\notin\operatorname{div}(i)$ for all $i\in I$. Using $J\subseteq I$ and by Lemma 5.2 we obtain that ${\boldsymbol{\mathit{Z}}}_{i}\vDash\mathrm{(div_{p,n})}$ for each $i\in I$ and ${\boldsymbol{\mathit{Z}}}_{j}\overrightarrow{\times}{\boldsymbol{\mathit{Z}}}_{0}\vDash\mathrm{(div_{p,n})}$ for each $j\in J$. Therefore, $\mathcal{V}_{I,J}\vDash\mathrm{(div_{p,n})}$ for each $p\notin I$.